OpenStudy (anonymous):

What is 0.0833333333333 as a fraction

OpenStudy (dan815):

8.333.. 8+1/3 25/3 1 ------- = ---------= --------= ----- 100 100 100 12

OpenStudy (anonymous):

i assume it is a non terminating decimal. let x=0.083333333333.... (1) 10 x=0.833333333333.......(2) 100 x=8.33333333333.......(3) 1000x=83.333333333.......(4) (4)-(3) gives 990 x=75 \[x=\frac{ 75 }{ 900 }=\frac{ 1 }{ 12 }\]

OpenStudy (anonymous):

correction write 900 in place of 990